OpenStudy (anonymous):

Solve the system. y = 4x + 12 and y = -2x + 6

OpenStudy (anonymous):

x=-1, y=8

Directrix (directrix):

Solving by substitution: (1) y = 4x + 12 (2) y = -2x + 6 ----------- in y = -2x + 6, substitute (4x + 12) for y in equation 2. This comes from the first equation. 4x + 12 = -2x +6 6x = -6 x=-1 From (1) y = 4x + 12, substitute -1 for x. y = 4 (-1) +12 = 8 Solution (-1,8)